Node Voltage Method Example 4: find v2 1,-24,1-3A, Ri-20,R,-40,R,-20,R,-30 . Label currents and voltages (polarities arbitrarily chosen) + R3- 1. 2. Choose Node d (v) as the reference node 3. Define remaining n -1 (3) voltages 2 (Vd 0) Va is independent V, is independent >Ve is independent 4. Apply KCL at nodes a, b, and c 1/27/2019 Node and Mesh Methods
Answer #1

At node 1

2= (Va-Vb)÷R1 +(Va-Vc)÷R3

Or 4= 2Va-Vb-Vc ------------1

At node 2

(Vb-Va)÷R1 +Vb÷R2 = 3

Or 3Vb-2Va=12-----------2

AT node 3

Vc÷R4 =3+(Vc-Va)÷R3

Or Vc-3Va =-9 -----------3

Putting value of Vc from eq 3 to eq 1 we get

4=-Va -Vb+9 or Vb-Va=5 ------4

Putting value of Vb-Va in eq 2 we get

Vb + 2×5 =12

Or Vb = 2 volt

Above is required value